The sum of two numbers is 148. The larger number is 16 more than twice the smaller number. What are the two numbers?
Answer by mananth(16949)

let one number be x
the other number will be 148-x
...
x=16+2(148-x)
x=16+296-2x
x+2x=312
3x=312
x= 104
other number = 148-104=44
...
104,44
...
m.ananth@hotmail.ca